# Analyticity of the Black Hole S-Matrix

### v1

@@ 1 Introduction
The analytic structure of the S-matrix encodes a great deal of information about the underlying dynamics of a quantum field theory (QFT). Basic principles such as causality, unitarity, and crossing symmetry impose stringent constraints on the location and nature of its singularities. For instance, poles identify the spectrum of intermediate states and reflect factorization properties, while discontinuities across branch cuts capture particle production and are controlled by unitarity Cutkosky:1960; Froissart:1961; Martin:1963UnitarityHE; Martin:1969ScatteringTheory; Eden:1966AnalyticSMatrix; Correia:2020xtr. Conversely, the existence of extended regions of analyticity can often be traced back to micro-causality StreaterWightman:1989PCT; Lehmann:1958Ellipse, the principle that observables at space-like separation must commute.
This understanding, which extends beyond perturbation theory, lies at the heart of the modern S-matrix bootstrap, which has been applied to a wide range of problems: from strongly-coupled gapped QFTs Paulos:2017fhb; Guerrieri:2018uew; Hebbar:2020ukp; Guerrieri:2023qbg; He:2023lyy; Cordova:2023wjp; Guerrieri:2024jkn; Guerrieri:2024ckc; Gumus:2024lmj; Cordova:2025bah; EliasMiro:2025rqo to mapping out the space of quantum UV completions of gravity Caron-Huot:2021rmr; Bellazzini:2015cra; Bern:2021ppb; Bellazzini:2021shn; Caron-Huot:2022ugt; Serra:2022pzl; Chiang:2022jep; Haring:2022cyf; Beadle:2025cdx; Chang:2025cxc and other effective field theories (EFTs) Bellazzini:2016xrt; Cheung:2016yqr; Bellazzini:2023nqj; Albert:2022oes; Serra:2022pzl. In this context, fundamental properties of an unknown ultraviolet (UV) theory can be translated into constraints on its low-energy EFT through dispersion relations Adams:2006sv; Arkani-Hamed:2020blm; deRham:2017avq; Bellazzini:2020cot; Sinha:2020win; Tolley:2020gtv; Caron-Huot:2020cmc; Beadle:2024hqg; Paulos:2017fhb; EliasMiro:2022xaa, allowing one to shape the space of Wilson coefficients that admit a consistent UV completion.
@@
where $x$ and $x^{\prime}$ denote the spatial coordinates of observer and source. This relation implies analyticity in the upper half of the complex frequency plane. However, this argument, often referred to as the signal model, does not immediately extend to the S-matrix (and indeed we show that it is violated). Concerning the lower half plane, it is well established that the retarded Green’s function showcases two types of singularities. The first are poles located symmetrically about the imaginary axis: the quasinormal modes Regge:1957td; Vishveshwara:1970cc; Chandrasekhar:1975zza; Berti:2009kk; Press:1971wr; Vishveshwara:1970zz; KonoplyaZhidenko2011; Ferrari:1984zz. These modes govern the exponential damping of the signal observed during the ringdown phase of gravitational-wave emission. The second set of singularities corresponds to a branch cut extending along the negative imaginary axis, known to control the late-time tail behavior of the signal Leaver:1986gd; price_nonspherical_1972_2439; price_nonspherical_1972_2419; Casals:2012ng; Casals:2015nja.
In contrast to the retarded Green’s function, the analytic structure of the transmission and reflection coefficients of the S-matrix has not yet been established (see work in this direction Castro:2013lba; Casals:2015nja; Motl:2003cd; Neitzke:2003mz). Naively, one might expect that the elastic amplitude $R(\omega)$ would inherit the analytic properties of the retarded Green’s function, as suggested by the asymptotic relation Andersson:1996cm
@@
Yet this expectation does not necessarily hold, due to the possible emergence of Stokes phenomena Stokes:1864; Barry:1989zz; Dingle:1973, which can arise when the source and the observer are taken to infinity, $x,x^{\prime}\to\infty$. [^1] The Stokes phenomenon Stokes:1864; Dingle:1973; Barry:1989zz can introduce singularities in the asymptotic limit of the Green’s function. In particular, because the wave equation has an irregular singular point at $x=\infty$, exponentially suppressed and enhanced terms can interchange dominance across the imaginary $\omega$ axis, leading to possible singularities of the reflection coefficient there. Stokes phenomena typically appear in asymptotic perturbative expansions and play a central role in the theory of resurgence Marino:2015yie; Aniceto:2018bis; Serone:2024uwz. In this work we fill this gap by establishing the analyticity properties of the transmission and reflection coefficients.
@@
We find that the analyticity guaranteed by causality for the retarded Green’s function in the UHP is not, by itself, sufficient to ensure analyticity of $R(\omega)$. The missing ingredient is a “locality” assumption on the potential: it admits an analytic continuation in the complex $x$-plane. This is typically verified for potentials that arise from a field theory, as is the case for the Schwarzschild gravitational potential that we study here.
@@ 2.2 Retarded Green’s function
The retarded causal condition (11) is established if $G_{R}(\omega,x,x^{\prime})$ is analytic for $\mathrm{Im}\,\omega>0$, which is in turn guaranteed if $\phi_{L}(\omega,x)$ and $\phi_{R}(\omega,x)$ are analytic for $\mathrm{Im}\,\omega>0$, and the Wronskian $W[\phi_{L},\phi_{R}](\omega)$ has no zeros for $\mathrm{Im}\,\omega>0$, i.e. no bound-states. [^3] While this would naively mean that the presence of bound states is incompatible with causality, it is important to note that if the potential $V(x)$ is bounded from below only a finite number of bound-states exists Yafaev:1992; Newton:1983JOT; Chadan:2002bound1D. In this case, usage of the Laplace transform instead of Fourier transform amounts to a deformation of the frequency plane contour above all the bound states which preserves the requirement of causality; refer to Appendix B for more details.
@@
These translate to causal boundary conditions for the retarded Green’s function, namely that an asymptotic observer at $x\to\pm\infty$ only sees waves propagating out from a source located at finite $x^{\prime}$. For example,
@@
Since the left outgoing solution $\phi_{L}(\omega,x)$ satisfies the same boundary condition as the physical solution $\phi(\omega,x)$ (6), they are related up to some normalization. This implies the asymptotic relation (2) between the reflection coefficient $R(\omega)$ and the retarded Green’s function $G_{R}(\omega,x,x^{\prime})$ when both the source and the observer are taken to infinity, $x,x^{\prime}\to\infty$.
@@ 2.3 Analyticity from causality
$\leq{1\over n!\,|\omega|^{n}}\left[\int_{-\infty}^{x}|V(x^{\prime})|\,dx^{\prime}\right]^{n}\,.\qquad\mathrm{Im}\,\omega\geq 0$ (24)
@@
Analyticity in UHP. Using these integral relations and the proven analyticty for $\chi_{L}(\omega,x)$, the analyticity of the connection coefficients follows. In particular, the integral representation for $A_{\text{in}}(\omega)$ (29) remains convergent given the bound (25), and moreover the proven analyticity of $\chi_{L}(\omega,x^{\prime})$ allows $A_{\text{in}}(\omega)$ to be analytically continued for $\text{Im}\,\omega>0$. [^5] This result also follows from the basic observation that $A_{\text{in}}(\omega)$ is given by the Wronskian (14) of two analytic functions.
@@
Analyticity for the reflection amplitude $R(\omega)$ requires additional assumptions on the potential, which we will now discuss.
@@ 2.4 Extended analyticity for the transmission amplitude
Moreover, we also assume that the potential remains bounded in this domain:
@@
and deform the contour of integration of 3, such that the new contour of integration is inside the domain (32).
@@
We can write the series expansion in (22) as
@@
$\leq{1\over n!\,|\omega|^{n}}\left[\int_{-\infty}^{y}|V(e^{i\theta}y^{\prime})|\,dy^{\prime}\right]^{n},\qquad\mathrm{Im}\,(e^{i\theta}\omega)>0.$ (40)
@@ 2.5 Extended analyticity for the reflection amplitude
Analyticity of $A_{\text{out}}(\omega)$. Proving analyticity of $A_{\text{out}}(\omega)$ requires a slightly more sophisticated argument. We start again with $\chi_{L}(\omega,x)$, but instead of slating the full contour over $x$ at a single angle $\theta$, we deform the integration from $\infty\to 0$ by an angle $\theta_{-}$ and from $0\to+\infty$ by a different angle $\theta_{+}$. We write this new contour $\gamma(x)$, shown in Figure 4, where
@@
depending on which branch $y$ and $y^{\prime}$ are on. Note that $y>y^{\prime}$ implies that a potential fourth piece proportional to $\Theta(-y)\Theta(y^{\prime})$ does not have support.
For $y<0$, the ordering of the integral (47) implies that all integration variables must be negative, $y_{j}<0$, meaning that there is no dependence on $\theta_{+}$. In this case we just recycle what we concluded in the previous analysis, namely the bound (41) replacing $\theta=\theta_{-}$.
@@ 2.6 Poles of the S-matrix: $A_{\text{in}}(\omega)=0$
In the above section, we have seen that the reflection amplitude is meromorphic in the blue and dotted region of the complex-$\omega$ plane shown in Figure 5. Within this region, there can be isolated poles arising from zeros of $A_{\text{in}}$. We find that they can only occur in the LHP and on the imaginary axis of the UHP.
@@
The relation (70) constrains the location of $\omega_{0}$ to one of two possibilities. Either $\omega_{0}$ is purely imaginary ($\mathrm{Re}\,\omega_{0}=0$), with both of sides of the equation vanishing. In this case, it is possible to have $\omega_{0}=i\kappa$, with $\kappa>0$, corresponding to a bound-state Mizera:2023tfe. [^9] While these bound-state poles are located in the UHP they do not contradict causality. In particular, using a Laplace transform instead of a Fourier transform in the definition of the retarded Green’s function (12) explicitly shows the preservation of causality. See Appendix B for details and examples.
@@
corresponding to a quasinormal mode (QNM). A QNM is a non-normalizable solution that diverges at spatial infinity and is exponentially damped in time. [^10] Note that the limit $b\to+\infty$, $a\to$-$\infty$ must be taken with care as both numerator and denominator in (71) become divergent. Their ratio should however give a finite result, given the left-hand-side, even though we have not shown this explicitly here. See Chapter 12 of Taylor:1972 for a different argument that does not suffer from this limitation. They are expected to play a role in the ringdown phase of black hole mergers Regge:1957td; Vishveshwara:1970cc; Chandrasekhar:1975zza; Berti:2009kk; Press:1971wr; Vishveshwara:1970zz; KonoplyaZhidenko2011; Ferrari:1984zz.
@@
Since $A_{\text{in}}(i\kappa)=0$, the Jost solution $\phi_{L}(i\kappa,x)$ is exponentially suppressed in both limits $x\to\pm\infty$, given the boundary conditions (15) and (26). If $V(x)>0$, each term on the LHS is nonnegative, enforcing $\phi_{L}(i\kappa,x)=0$. Therefore, all bound-state solutions are trivial and no bound-state poles are present for positive potentials.
@@ 3.1 Analyticity domains of the S-matrix
The Pöschl-Teller potential (76) is bounded, satisfying (17), and decays exponentially,
@@
Moreover, $V(x)$ can be analytically continued into the complex-$x$ plane, satisfying conditions (32) and (33) with $\theta_{\text{max}}=\pi/2$. The potential is plotted in Figure 7, both for real $x$ and in the complex-$x$ plane.
@@
Figure 7: The Pöschl–Teller potential barrier $V(x)$ with $\alpha=1,\lambda=1/2+0.3\,i$ for Left: $x\in\mathbb{R}$ and Right: $x\in\mathbb{C}$. The height of the complex plot is the magnitude of the potential $|V(x)|$, while the color corresponds to the argument of the potential $\arg V(x)$. The potential has an infinite number of poles on the imaginary axis, located at $x=\tfrac{i}{\alpha}\left(\tfrac{\pi}{2}+2\pi n\right)$, $n\in\mathbb{Z}$, and decays exponentially on the real axis as $x\rightarrow\pm\infty$. The solid line on the right plot shows the integration contour for real $x$, and the dashed line shows the integration contour for $x=e^{i\theta}y$.
With these conditions met, following the analysis of Section 2.4, we also show that the transmission and reflections coefficients, $T(\omega)$ and $R(\omega)$, have analyticity domains given in (74) and (75), respectively. In particular, for $T(\omega)$ we show that it is analytic in the UHP and meromorphic in the LHP except on the negative imaginary axis. For the reflection amplitude $R(\omega)$ we show that it shares the same analytic structure of $T(\omega)$ in the LHP, and in the UHP it is analytic except along the imaginary axis, where it could have singularities.
In Appendix A, we show that for exponentially decaying potentials, the analytic domain extends to an additional horizontal strip whose width is set by the exponential falloff (see Figure 6). For the Pöschl–Teller potential, this implies extra analyticity in the region $\alpha<\text{Im}\,\omega<\alpha$.

### v2

@@ 1 Introduction
The analytic structure of the S-matrix encodes a great deal of information about the underlying dynamics of a quantum field theory (QFT). Basic principles such as causality, unitarity, and crossing symmetry impose stringent constraints on the location and nature of its singularities. For instance, poles identify the spectrum of intermediate states and reflect factorization properties, while discontinuities across branch cuts capture particle production and are controlled by unitarity Cutkosky:1960; Froissart:1961; Martin:1963UnitarityHE; Martin:1969ScatteringTheory; Eden:1966AnalyticSMatrix; Correia:2020xtr. Conversely, the existence of extended regions of analyticity can often be traced back to micro-causality StreaterWightman:1989PCT; Lehmann:1958Ellipse, the principle that operators at space-like separation must commute.
This understanding, which extends beyond perturbation theory, lies at the heart of the modern S-matrix bootstrap, which has been applied to a wide range of problems: from strongly-coupled gapped QFTs Paulos:2017fhb; Guerrieri:2018uew; Hebbar:2020ukp; Guerrieri:2023qbg; He:2023lyy; Cordova:2023wjp; Guerrieri:2024jkn; Guerrieri:2024ckc; Gumus:2024lmj; Correia:2025uvc; Cordova:2025bah; EliasMiro:2025rqo to mapping out the space of quantum UV completions of gravity Caron-Huot:2021rmr; Bellazzini:2015cra; Bern:2021ppb; Bellazzini:2021shn; Caron-Huot:2022ugt; Serra:2022pzl; Chiang:2022jep; Haring:2022cyf; Beadle:2025cdx; Chang:2025cxc and other effective field theories (EFTs) Bellazzini:2016xrt; Cheung:2016yqr; Bellazzini:2023nqj; Albert:2022oes; Serra:2022pzl. In this context, fundamental properties of an unknown ultraviolet (UV) theory can be translated into constraints on its low-energy EFT through dispersion relations Adams:2006sv; Arkani-Hamed:2020blm; deRham:2017avq; Bellazzini:2020cot; Sinha:2020win; Tolley:2020gtv; Caron-Huot:2020cmc; Beadle:2024hqg; Paulos:2017fhb; EliasMiro:2022xaa, allowing one to shape the space of Wilson coefficients that admit a consistent UV completion.
@@
where $x$ and $x^{\prime}$ denote the spatial coordinates of observer and source. This relation implies analyticity in the upper half of the complex frequency plane. However, this argument, often referred to as the signal model Camanho:2014apa, does not immediately extend to the S-matrix (and indeed we show in this work that it is violated).
Concerning the lower half plane, it is well established that the retarded Green’s function showcases two types of singularities. The first are poles located symmetrically about the imaginary axis: the quasinormal modes Regge:1957td; Vishveshwara:1970cc; Chandrasekhar:1975zza; Berti:2009kk; Press:1971wr; Vishveshwara:1970zz; KonoplyaZhidenko2011; Ferrari:1984zz. These modes govern the exponential damping of the signal observed during the ringdown phase of gravitational-wave emission Berti:2009kk. The second set of singularities corresponds to a branch cut extending along the negative imaginary axis, known to control the late-time tail behavior of the signal Leaver:1986gd; price_nonspherical_1972_2439; price_nonspherical_1972_2419; Casals:2012ng; Casals:2015nja.
In contrast to the retarded Green’s function, the analytic structure of the transmission and reflection coefficients of the black hole S-matrix has not yet been established (see Castro:2013lba; Casals:2015nja; Motl:2003cd; Neitzke:2003mz; Arnaudo:2025uos for work in this direction). Naively, one might expect that the elastic amplitude $R(\omega)$ would inherit the analytic properties of the retarded Green’s function, as suggested by the asymptotic relation Andersson:1996cm
@@
Yet, as we point out in this work, this expectation does not necessarily hold, due to the possible emergence of Stokes phenomena Stokes:1864; Barry:1989zz; Dingle:1973, which can arise when the source and the observer are taken to infinity, $x,x^{\prime}\to\infty$. [^1] The Stokes phenomenon Stokes:1864; Dingle:1973; Barry:1989zz can introduce singularities in the asymptotic limit of the Green’s function. In particular, because the wave equation has an irregular singular point at $x=\infty$, exponentially suppressed and enhanced terms can interchange dominance across the imaginary $\omega$ axis, leading to possible singularities of the reflection coefficient there. Stokes phenomena typically appear in asymptotic perturbative expansions and play a central role in the theory of resurgence Marino:2015yie; Aniceto:2018bis; Serone:2024uwz. In this work we fill this gap by establishing the analyticity properties of the transmission and reflection coefficients.
@@
We find that the analyticity guaranteed by causality for the retarded Green’s function in the UHP is not, by itself, sufficient to ensure analyticity of $R(\omega)$. The missing ingredient is a “locality” assumption on the potential: that it admits an analytic continuation in the complex $x$-plane. This is typically verified for potentials that arise from a field theory, as is the case for the Schwarzschild gravitational potential that we study here.
@@ 2.2 Retarded Green’s function
The retarded causal condition (11) is established if $G_{R}(\omega,x,x^{\prime})$ is analytic for $\mathrm{Im}\,\omega>0$, which is in turn guaranteed if $\phi_{L}(\omega,x)$ and $\phi_{R}(\omega,x)$ are analytic for $\mathrm{Im}\,\omega>0$, and the Wronskian $W[\phi_{L},\phi_{R}](\omega)$ has no zeros for $\mathrm{Im}\,\omega>0$. [^3] These Wronskian zeros would correspond to poles in the UHP of the retarded Green’s function, typically associated with bound-states. While their presence would naively be incompatible with causality, it is important to note that if the potential $V(x)$ is bounded from below only a finite number of bound-states exists Yafaev:1992; Newton:1983JOT; Chadan:2002bound1D. In this case, usage of the Laplace transform instead of Fourier transform amounts to a deformation of the frequency plane contour above all the bound states which preserves the requirement of causality; refer to Appendix B for more details.
@@
These translate to causal boundary conditions for the retarded Green’s function, namely that an asymptotic observer at $x\to\pm\infty$ only sees waves propagating outwards from a source located at finite $x^{\prime}$. For example,
@@
Since the left outgoing solution $\phi_{L}(\omega,x)$ satisfies the same boundary condition as the physical solution $\phi(\omega,x)$ (6), they are related up to some normalization. This leads to the asymptotic relation (2) between the reflection coefficient $R(\omega)$ and the retarded Green’s function $G_{R}(\omega,x,x^{\prime})$ when both the source and the observer are taken to infinity, $x,x^{\prime}\to\infty$.
@@ 2.3 Analyticity from causality
$={1\over n!\,|\omega|^{n}}\left[\int_{-\infty}^{x}|V(x^{\prime})|\,dx^{\prime}\right]^{n}\,.\qquad\mathrm{Im}\,\omega\geq 0$ (24)
@@
Analyticity in UHP. Using these integral relations and the proven analyticty for $\chi_{L}(\omega,x)$, some analyticity domain for the connection coefficients follows. In particular, the integral representation for $A_{\text{in}}(\omega)$ (29) remains convergent given the bound (25), and moreover the proven analyticity of $\chi_{L}(\omega,x^{\prime})$ allows $A_{\text{in}}(\omega)$ to be analytically continued for $\text{Im}\,\omega>0$. [^5] This result also follows from the basic observation that $A_{\text{in}}(\omega)$ is given by the Wronskian (14) of two analytic functions.
@@
Analyticity in the lower-half-plane of $T(\omega)$ requires additional assumptions on the potential, which we will now discuss.
@@ 2.4 Extended analyticity for the transmission amplitude
Moreover, we also assume that the potential decays sufficiently fast in this wedge:
@@
and deform the contour of integration of each Born iteration (22), such that the new contour of integration is inside the domain (32), as shown in Figure 3.
@@
We can write the series expansion in (22) along the deformed contour as
@@
$={1\over n!\,|\omega|^{n}}\left[\int_{-\infty}^{y}|V(e^{i\theta}y^{\prime})|\,dy^{\prime}\right]^{n},\qquad\mathrm{Im}\,(e^{i\theta}\omega)>0.$ (40)
@@ 2.5 Extended analyticity for the reflection amplitude
We now show that a similar argument involving contour deformation into the complex $x$-plane allows to prove an extended domain of analyticity for the reflection coefficient $R(\omega)$. Given the relation (28) and having established analyticity for $A_{\text{in}}(\omega)$ it remains to show analyticity for $A_{\text{out}}(\omega)$.
Analyticity of $A_{\text{out}}(\omega)$. We start again with extending the analyticity domain of $\chi_{L}(\omega,x)$, but instead of slating the full contour over $x$ at a single angle $\theta$, we deform the integration from $\infty\to 0$ by an angle $\theta_{-}$ and from $0\to+\infty$ by a different angle $\theta_{+}$. We write this new contour $\gamma(x)$, shown in Figure 4, where
@@
depending on which branch $y$ and $y^{\prime}$ are on. Note that $y>y^{\prime}$ implies that a possible fourth piece proportional to $\Theta(-y)\Theta(y^{\prime})$ does not have support.
Notice also that for $y<0$, the ordering of the integral (47) implies that all integration variables must be negative, $y_{j}<0$, meaning that there is no dependence on $\theta_{+}$. In this case we just recycle what we concluded in the previous analysis, namely the bound (41) replacing $\theta=\theta_{-}$.
@@ 2.6 Poles of the S-matrix: $A_{\text{in}}(\omega)=0$
In the above section, we have seen that the reflection amplitude is meromorphic in the blue and dotted region of the complex-$\omega$ plane shown in Figure 5. Within this region, there can be isolated poles arising from zeros of $A_{\text{in}}$. We now show that they can only occur either everywhere in the LHP or along the imaginary axis of the UHP.
@@
The relation (70) constrains the location of $\omega_{0}$ to one of two possibilities. Either $\omega_{0}$ is purely imaginary ($\mathrm{Re}\,\omega_{0}=0$), with both of sides of the equation vanishing. In this case, it is possible to have $\omega_{0}=i\kappa$, with $\kappa>0$, corresponding to a bound-state Mizera:2023tfe. [^9] While these bound-state poles are located in the UHP they do not contradict causality. In particular, using a Laplace transform instead of a Fourier transform in the definition of the retarded Green’s function (12) explicitly shows the preservation of causality. See Appendix B for details.
@@
corresponding to a quasinormal mode (QNM). A QNM is a non-normalizable solution that diverges at spatial infinity and is exponentially damped in time. [^10] Note that in this case the limit $b\to+\infty$, $a\to$-$\infty$ must be taken with care as both numerator and denominator in (71) become divergent. Their ratio should however give a finite result given that the left-hand-side of (71) is finite, even though we have not shown this explicitly here. See Chapter 12 of Taylor:1972 for a different argument that does not suffer from this limitation. They are expected to play a role in the ringdown phase of black hole mergers Regge:1957td; Vishveshwara:1970cc; Chandrasekhar:1975zza; Berti:2009kk; Press:1971wr; Vishveshwara:1970zz; KonoplyaZhidenko2011; Ferrari:1984zz.
@@
Since $A_{\text{in}}(i\kappa)=0$, the Jost solution $\phi_{L}(i\kappa,x)$ is exponentially suppressed in both limits $x\to\pm\infty$, given the boundary conditions (15) and (26), setting the RHS of (73) to zero. If $V(x)>0$, each term on the LHS is nonnegative, enforcing $\phi_{L}(i\kappa,x)=0$. Therefore, all bound-state solutions are trivial and no bound-state poles are present for positive potentials.
@@ 3.1 Analyticity domains of the S-matrix
The Pöschl-Teller potential (76) is bounded, satisfying (17), and decays exponentially
@@
Moreover, $V(x)$ can be analytically continued into the complex-$x$ plane, and has poles on the imaginary-$x$ axis. Therefore, it satisfies conditions (32) and (33) with $\theta_{\text{max}}=\pi/2-\epsilon$ with $\epsilon>0$ arbitrarily small. The potential is plotted in Figure 7, both for real $x$ and in the complex-$x$ plane.
@@
Figure 7: The Pöschl–Teller potential barrier $V(x)$ with $\alpha=1,\lambda=1/2+0.3\,i$ for Left: $x\in\mathbb{R}$ and Right: $x\in\mathbb{C}$. The height of the complex plot is the magnitude of the potential $|V(x)|$, while the color corresponds to the argument of the potential $\arg V(x)$. The potential has an infinite number of poles on the imaginary axis, located at $x=\tfrac{i}{\alpha}\left(\tfrac{\pi}{2}+2\pi n\right)$, $n\in\mathbb{Z}$, and decays exponentially on the real axis as $x\rightarrow\pm\infty$. The solid line on the right plot shows the integration contour for real $x$, and the dashed line shows the integration contour for $x=e^{i\theta}y$. We can take $\theta\to\pi/2-\epsilon$ with $\epsilon>0$ arbitrarily small without encountering singularities.
With these conditions met, following the analysis of Section 2.4, we conclude that the transmission and reflections coefficients, $T(\omega)$ and $R(\omega)$, have analyticity domains given in (74) and (75), respectively. In particular, for $T(\omega)$ we show that it is analytic in the UHP and meromorphic in the LHP except on the negative imaginary axis. For the reflection amplitude $R(\omega)$ we show that it shares the same analytic structure of $T(\omega)$ in the LHP, and in the UHP it is analytic except along the imaginary axis, where it could have singularities.
In Appendix A, we also showed that for exponentially decaying potentials, the analytic domain in $\omega$ extends to an additional horizontal strip whose width is set by the exponential falloff (see Figure 6). For the Pöschl–Teller potential, this implies extra analyticity in the region $\alpha<\text{Im}\,\omega<\alpha$.